OpenStudy (anonymous):

what's the inverse of f(x)=|x+3| ?? x is > or = to -3

OpenStudy (watchmath):

Since \(x\geq -3\) then \(|x+3|=x+3\). For \(y=x+3\) then \(x=y-3\). So \(f^{-1}(x)=x-3\)

OpenStudy (anonymous):

Thank, I'm always nervous about absolute value

OpenStudy (anonymous):

Thank you*

OpenStudy (anonymous):

watchmath got a second?

OpenStudy (watchmath):

sure, where do you want to talk?

OpenStudy (anonymous):

i want you to check a rather easy probability problem i just did which doesn't correspond to any of the multiple choice anwers the student gave.